[Brahms-l] Fw: A-SPY news: last time meeting June 21, 2005 ... schedule changes ...

From: flemming videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Mon Jun 20 2005 - 16:43:45 EDT
----------------------------------------------------------------
Flemming Videbaek
Physics Department
Brookhaven National Laboratory

e-mail: videbaek@bnl.gov
phone: 631-344-4106
----- Original Message ----- 
From: Kin Yip 
To: kinyip@fnal.gov 
Sent: Monday, June 20, 2005 4:37 PM
Subject: A-SPY news: last time meeting June 21, 2005 ... schedule changes ...


====== Acting Scheduling Physicist -YIP (ie. A-SPY) news channel ========

***********************************************************************************************

(0) The last Time Meeting of this run will be tomorrow (June 21, 2005) at 1:30 pm at the Snyder 
     Room.

***********************************************************************************************

(1) Schedule changes...

--- There will be machine development of 3 hours-ish this Tuesday (orginally was today but it's 
    now changed), after the 1st store of the day ends.

--- The last Accelerator Physics Expt. section will now be from 09:00 (am) June 24, 2005  
                                                                                     to    01:00 (am) June 25, 2005.
    The 01:00 on June 25, 2005 is absolute !   No more change.

--- The "AGS cold snake" work would hopefully stop around the above time.   Preparation for the 
     NASA AGS run would hopefully start on Sat. (fixing the water hose at the F10 magnet 
     in AGS and the power supply test etc.).   And hopefully, the NASA physics experimenters
     can have beams to use not too late next week ..........

***********************************************************************************************

(2) Alarm fix/Transverse polarization around last Thursday 

The run with transverse polarization ended around 10:00 am last Thursday (June 16) and
PHENIX transverse seems to have gotten 163 nb-1 (more than 100 nb-1 they'd originally
asked for).

The alarm fixes at 12 o'clock and 7 o'clock were within 2 hours --- the original optimistic 
estimate.   They have done a lot of mopping to dry the floor and replaced the two detectors
which gave problems.

Two "SPY news" readers have sent me email (after reading my last "SPY news") to advise 
us to use heater instead of AC/dehumidifier.    At the end, we have kind of switched both 
heater and dehumidifier at 7 o'clock.  Heater is to raise the dew-point for the air coming in 
(so that there will be less condensates on the electrical equipment) and dehumidifer 
intends to suck any moisture in the tunnel.

***********************************************************************************************

(3) The expt. at the Radiation Effect Facility (REF) has just started ...

I have seen the Radiation Safety Committee Chairman signs off the last part of the
REF check-off list.  So, apparently, they can finally start to do their laser stripping
expt. (to strip electrons from H-) ... even though there are not too many days left.

***********************************************************************************************

======= ====== ======= ======= ====== ======= ======= ====== =======


Kin Yip
Micro-Management Ltd.

Tel.:    x4116
Pager: x4300 

=================================================================

Disclaimer: Whatever is being said here is not always 100% accurate 
though I try my best to avoid being 100% inaccurate :-) 

+++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++





_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jun 20 16:41:51 2005

This archive was generated by hypermail 2.1.8 : Mon Jun 20 2005 - 16:41:56 EDT